Weather Report for September 18, 2015

Tonight was the first night of the campaign that the dome could be kept open all night long. It was not without weather problems as cirrus was in the sky to various degrees all night long. The clouds were worse after midnight and no photometry was attempted. Photometry in the early part of the night will have to be verified to see if thin cirrus invalidated the measurements. Spectropolarimetry continued throughout the night and 20 blazars were observed. The seeing was generally quite good, being 1-1.5 arcsec.